The sum of twice a number and 4 less than the number is the same as the diffrence between -32 and the number. What is the number?
Answer by ewatrrr(24785)


Hi,
Let x represent the value of the number
Write as we Read***
2x + (x-4) = x - (-32)
3x - 4 = x + 32
2x = 36
x = 18
checking our answer
36 + 14 = 18 + 32 = 50